When one molecule of methane combusts with oxygen, 5.74 X 10-21 J is released as heat. How many molecules of methane must undergo combustion to heat up 1 gram of water by 1 degree C. note that this process requires 4.184 J. IF one molecule of methane has a mass of 2.66 X 10-23g, how many total grams of methane are required?
Answer #1

Answer

0.01939g

Explanation

Heat released by one methane molecule = 5.74×10-21J

Heat required to raise 1g of water to 1℃ = 4.184J

Number of methane molecule required =

4.184J / 5.74 ×10-21J = 7.289× 1020molecules

mass of 1 molecule of methane = 2.66 ×10-23g

mass of 7.289 × 1020molecules = 7.289 ×1020molecule × 2.66 ×10-23g/molecule = 0.01939g